# ViViDex: Learning Vision-based Dexterous Manipulation from Human Videos

### v1

@@ IV-D Evaluation of our policy on Pour and Place inside Tasks
We further evaluate our approach on the pour and place inside tasks. Table VII shows the quantitative results. Rows S1 to S4 are state-based policies reported in [23] where they use 91 videos and 101 videos for pour and place inside tasks respectively. While we only adopt one video in training, our state-based policy S7 achieves 97% and 68% success rates on the two tasks respectively, significantly outperforming 27% and 31% success rates using their best models. Benefiting from high-quality trajectories generated by our state-based policies, our visual policies V5 also achieve good performance on these two tasks. Figure 6 illustrates qualitative examples from our visual policy for these two tasks.
@@ V Conclusion
We introduce ViViDex, a new framework to learn vision-based dexterous manipulation from human videos. To this end, we first extract reference hand and object trajectories from human videos and train state-based policies via RL to generate diverse physically plausible trajectories as the training data for the vision-based policy. Then, the visual policy takes robot proprioceptive states and 3D point clouds as inputs and further transforms 3D point clouds into different coordinate systems to achieve better performance. We conduct extensive ablations experiments to validate the effectiveness of our ViViDex system. We show that our visual policies can outperform state-of-the-art accuracy by a great margin on three dexterous manipulation tasks. Future work aims at leveraging internet videos for acquiring more general dexterous manipulation skills and investigating advanced 3D pose estimators.
Acknowledgement: This work was granted access to the HPC resources of IDRIS under the allocation AD011013147 made by GENCI. It was funded in part by the French government under management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR19-P3IA-0001 (PRAIRIE 3IA Institute) and the ANR project VideoPredict (ANR-21-FAI1-0002-01).

### v2

@@ IV-D Evaluation of our policy on Pour and Place inside Tasks
We further evaluate our approach on the pour and place inside tasks and report results in Table V. Rows L1-L4 show results for state-based policies [24] using the Adroit hand as well as 91 and 101 video demonstrations for pour and place inside tasks respectively. While we use only single human video for training, our state-based policy L5 for the Adroit hand achieves 97% and 68% success rates on these two tasks respectively, significantly outperforming 27% and 31% success rates corresponding to the best models of [24]. By taking advantage of high-quality trajectories generated by L5, our visual policy L6 also achieves high performance.
TABLE VI: Success rate of visual policies on relocate task using the real robot. We evaluate 10 episodes for each object.
Methods Unified policy Seen Unseen Avg.
Mb. Tc. Sb. Lc. Mug Avg.
R1 BC $\times$ 10/10 8/10 9/10 8/10 9/10 0.88 -
R2 BC 9/10 7/10 7/10 5/10 8/10 0.72 0.58
R3 Diffusion 10/10 7/10 8/10 7/10 8/10 0.80 0.68
Figure 4: Illustrations of our real-world robot experimental setup and the performance of our proposed ViViDex algorithm.
IV-E Real world experiments
To further demonstrate the advantages of ViViDex, we evaluate its performance on real-world dexterous manipulation. As shown in Figure 4(a), we use a UR5 robotic arm equipped with an Allegro hand and a single RealSense D435 RGB-D camera. Since point clouds for the real-world experiments are noisier than those in simulation, we use the state-based policy to collect data in the real world and then learn a unified visual policy based on real data. We run state-based policies in the simulator for specific object locations and run them on the real robot by placing the object in the same locations. This allows us to collect 3D scene point clouds and robot states for each execution step and build the real-robot training dataset. We collect five robot trajectories for each object and train the visual policy using our collected real-robot data. As shown in Figure 4(b), our visual policy can then manipulate the object. We summarize the quantitative results on the real robot in Table VI. We adopt the initial pose for seen objects from the first row of Figure 2 and include five unseen objects for evaluation: cracker box, spray bottle, bleach cleanser, water bottle and pudding box. R1 trains five policies separately and achieve an average success rate of 88%. R2 and R3 learn unified policies for five objects and demonstrate the ability to grasp unseen objects. The diffusion policy R3 also achieves better performance than BC in real experiments.
@@ V Conclusion
We introduce ViViDex, a new framework for learning vision-based dexterous manipulation from human videos. Our approach extracts reference trajectories from human videos and uses them as a reward in training state-based policies with RL. This generates diverse physically plausible trajectories. We rollout these trajectories for training a visual policy, which takes as inputs robot proprioceptive states and 3D point clouds. To enhence visual policies, we further transforms 3D point clouds into hand-centered coordinate systems. We conduct extensive ablation experiments to validate the effectiveness of ViViDex on different simulators and a real robot. We show that our visual policies outperform state-of-the-art accuracy on different tasks by a significant margin and generalize to unseen objects. Future work aims at leveraging internet videos for acquiring more general dexterous manipulation skills and investigating advanced 3D pose estimation algorithms.
Acknowledgement: We thank Ricardo Garcia Pinel for help with the camera calibration, Yuzhe Qin for clarifications about DexMV and suggestions on the real robot depolyment. This work was granted access to the HPC resources of IDRIS under the allocation AD011013147 made by GENCI. It was funded in part by the French government under management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR19-P3IA-0001 (PRAIRIE 3IA Institute) and the ANR project VideoPredict (ANR-21-FAI1-0002-01).